OpenStudy (anonymous):

simplify the qoutient e^4/e^3

OpenStudy (anonymous):

Consider this: a^n/a^z=a^(n-z)

OpenStudy (anonymous):

so in your case you would have: e^4/e^3=e^(4-3)=e^1=e

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

e